OpenStudy (kanard):

Question 9:

OpenStudy (kanard):

@jhonyy9

OpenStudy (kanard):

@TheSmartOne

jhonyy9 (jhonyy9):

so just use the rule of 3 simple 1 sec produce 3,9 *10^33 ergs of... 2,45*10^3 sec ----- x ergs ... ------------------------------ x = 2,45 *10^3 *3,9*10^33 / 1 = ? 2,45*10^? so this number not is sure that i have saw right - how many is there ?

OpenStudy (kanard):

It is 2.45*10^5

OpenStudy (kanard):

I did that but it just comes out to this 9.555e+38

jhonyy9 (jhonyy9):

ok so than you need multiplie x = 2,45*10^5 *3,9 *10^33 = ?